# SEFL: Enhancing Educational Assignment Feedback with LLM Agents

### v2

@@ Abstract
Providing high-quality feedback to student assignments is crucial for student success, but it is constrained by time and costs. In this work, we introduce Synthetic Educational Feedback Loops (SEFL), a synthetic data framework designed to generate data that resembles immediate, on-demand feedback at scale without relying on extensive, real-world student assignments. To get this type of data, two large language models (LLMs) operate in teacher–student roles to simulate assignment completion and formative feedback, generating synthetic pairs of student work and corresponding critiques and actionable improvements from a teacher. With this data, we fine-tune smaller, more computationally efficient LLMs on these synthetic pairs, enabling them to replicate key features of high-quality, goal-oriented feedback. Unlike personalized tutoring approaches that offer multi-turn, individualized instruction, SEFL specifically focuses on replicating the teacherstudent assignment feedback loop in higher education. Through comprehensive evaluations with four LLM judges and three human experts, we demonstrate that SEFL-tuned models outperform both their non-tuned counterparts in feedback quality and an existing baseline. The potential for societal impact is reinforced by extensive qualitative comments by ratings by human stakeholders—both students and higher education instructors. All in all, SEFL has substantial potential to transform feedback processes for higher education and beyond.
Code — https://github.com/jjzha/sefl
Datasets and Models — https://huggingface.co/collections/jjzha/sefl-synthetic-educational-feedback-loops-67b48768dab5123a2e3e9d69
1 Introduction
@@
Constructive feedback is a cornerstone of higher education, promoting critical thinking and fostering deeper understanding (Hattie 2008; Costello and Crane 2013). In higher education settings, however, providing consistent, high-quality feedback is complicated by privacy, consent, and transparency considerations in data collection (Fischer et al. 2020; Suresh et al. 2022; Demszky and Hill 2023; Wang and Demszky 2024; Wang et al. 2024b; Lindsay et al. 2024), in addition to being a labor-intensive task. Advances in language technology offer opportunities to automate and augment feedback processes, addressing these limitations.
@@
In particular, LLMs have shown progress in education (Wang et al. 2024c), including automated grading (Ke and Ng 2019; Ramesh and Sanampudi 2022; Stahl et al. 2024) and personalized tutoring (Yun et al. 2024; Liu et al. 2024c; Rooein and Hovy 2024; Ross and Andreas 2024; Kwon et al. 2024; Zhang et al. 2024, 2025; Wang et al. 2024a). Yet, automating teacher–student assignment feedback with LLMs remains an open question. We seek to answer: RQ. How can synthetic teacher–student interactions generated by LLMs be leveraged to enable scalable and effective educational student assessment feedback?
@@
Here, we introduce Synthetic Educational Feedback Loops (SEFL), a framework that generates synthetic teacher–student interactions using LLM agents. In this framework, two LLMs, one acting as the teacher and the other as the student, simulating formative feedback workflows (Conole and Oliver 2006; Nicol 2007). To induce better feedback from LLMs this synthetic data from the agents is then used to fine-tune smaller autoregressive language models, allowing the development of scalable educational feedback systems that can operate efficiently on modest computational infrastructure, available in higher education institutions without the need for privacy sensitive data.
@@
We present several findings. (i) Through empirical and qualitative analysis of expert annotator comments, we find that larger models tend to provide more actionable, goal-oriented, and user-friendly feedback. They are also more consistent and better at supporting student autonomy, showing the potential for use in real-world contexts. (ii) Empirically, SEFL-tuned models outperform their non-tuned versions in win rate evaluations by 4 LLM judges and 3 human experts in giving assessment feedback. (iii) We observe strong agreement among human evaluators regarding feedback quality. (iv) We compare our approach to Book2Dial (Wang et al. 2024a), showing that SEFL achieves a higher win rate according to 3 of the 4 LLM judges.
@@
We contribute the following in this work:
SEFL: A framework that simulates teacher student feedback loops with paired language model agents.
19,841 assignment–feedback pairs generated by SEFL to fine-tune smaller language models.
A comprehensive mix of human and language model evaluation that highlights strengths and limitations of SEFL with extensive qualitative analysis of the feedback provided by human experts.
An open source release of all models, code, and data.
2 Related Work
@@
Large language models are now supporting a broad spectrum of educational tasks. In automated grading, they score short answers, essays, and even programming assignments with accuracy that approaches expert instructors, which eases faculty workload and releases time for mentoring (Ke and Ng 2019; Ramesh and Sanampudi 2022; Stahl et al. 2024). For personalized tutoring, conversational agents powered by these models adapt explanations, hints, and examples to each learner’s background knowledge and preferred style, producing measurable gains in engagement and achievement (Yun et al. 2024; Liu et al. 2024c; Rooein and Hovy 2024; Ross and Andreas 2024; Kwon et al. 2024; Zhang et al. 2024). Research on peer learning shows that the same technology can mediate small-group discussions, suggest prompts, and highlight diverse viewpoints, leading to richer collaboration (Bauer et al. 2023). In mathematics, aligning word problems and proofs to grade-level objectives has been automated with encouraging results, reducing the manual effort needed to curate question banks (Botelho et al. 2023). Critical thinking curricula benefit as well; LLMs can challenge students to justify claims, detect fallacies, and refine arguments in real time (Guerraoui et al. 2023). The models have also begun to assist scholars: studies report successful use for screening literature, summarizing drafts, and aligning reviewer comments with revision plans (Liang et al. 2024; Sonkar et al. 2024). Complementing these functions are analytics tools that track learning trajectories and surface early warnings when a student slips behind (Schwarz et al. 2018; Aslan et al. 2019; Alrajhi et al. 2021).
Despite this growing body of work, prior studies have rarely targeted the systematic generation of rich feedback on open-ended student submissions and assignments at scale. The present study addresses that gap by using LLMs to create extensive comment sets that teachers can accept as-is or trim to taste. Decades of scholarship define effective feedback as goal-oriented, actionable, timely, user-friendly, and consistent while fostering self-evaluation (Carless et al. 2011; Wiggins 2012). Feedback that rambles can confuse learners, so concise wording is usually preferable, and comments that arrive soon after the original effort drive steady improvement (Wiggins 2012). By producing immediate responses and tailoring suggestions to rubric criteria, LLM based systems stand well positioned to satisfy these guidelines while operating at classroom and institution scale.
@@
Recent research shows how collaborative agentic LLMs can synthesize large-scale interactional datasets for educational tasks. For example, CAMEL (Li et al. 2023) uses cooperative role-based dialogues to achieve shared objectives, while SimSeek (Kim et al. 2022) uses agent-based conversations to build comprehensive information-seeking datasets. In education, SocraticLM (Liu et al. 2024b) simulates Socratic tutoring through multi-turn dialogue, and Book2Dial (Wang et al. 2024a) generates teacher-student conversations from textbooks. In contrast, SEFL focuses on concise teacher-student feedback loops rather than extended instructional dialogues. While Nair et al. (2024) explore iterative revisions, SEFL generates diverse feedback pairs from assignment-answer-feedback tuples, enabling fine-tuning of smaller, cost-effective models for large-scale use.
3 Synthetic Educational Feedback Loops
3.1 Synthetic Data Generation
@@
We use a two-agent framework (Wu et al. 2023) to simulate a feedback loop as in higher education. Both the teacher and student roles are simulated by two separate Llama-3.1-70B models for a two-turn conversation.111Note that if we mention a model, it is always the post-trained version (i.e., -Instruct). The models are tasked to generate assignmentanswerfeedback tuples. First, the student-agent asks for an assignment using Fineweb-Edu (Lozhkov et al. 2024) texts (Figure 1), known for its educational content based on LLM judgments. Second, the teacher-agent creates an assignment that can be of any domain, e.g., math, humanities, role-playing. Then, the student-agent ( ) submits assignments containing a number of explicit errors, and the teacher-agent ( ) provides feedback. We investigate both Qwen2.5-72B and Llama-3.1-70B for interactions. First, we generate 5,000 interaction tuples with each model, where we validate the output as an initial step to investigate initial feedback quality.
@@
We show in Table 1 the results of this first experiment. Out of 5,000 generated examples, Llama-3.1-70B generates 2,513 valid examples (i.e., valid JSON format and each feedback point refers to an error) compared to Qwen2.5-72B with 454 valid examples. For a further check, we use BERTScore (Zhang et al. 2020) as a proxy to see whether each error-feedback pair of the valid generations relate to each other.222We only calculate it of the samples where both error and feedback have the same number of generations. We show that, regardless of Llama-3.1-70B generating more valid examples, the BERTScore (0.877) stays in a similar range as Qwen2.5-72B (0.919); in both cases indicating high similarity. As a last qualitative check, we experimented with several prompts and consolidated into a final prompt found in the supplementary material. Finally, we use Llama-3.1-70B-generated data as the basis for all subsequent model fine-tuning as it generated more valid examples.
@@
Statistics.
After the 5K generated examples, we continue generating example pairs and end up with 19.8K teacher–student feedback pairs. In Table 2, we present the final dataset statistics. We highlight that the generation lengths for each agent are intentionally kept concise (170 subword tokens), based on the hypothesis that overly lengthy feedback may be counterproductive, especially for short assignments. This is in line with observations from Ferguson (2011), who finds that students tend to favor brief comments. We argue that balancing supportive and critical feedback is crucial as, by default, LLMs often produce excessively verbose responses, which can influence the preferences of both humans and language models (Saito et al. 2023).
4 Methodology
4.1 Fine-Tuning Large Language Models
We divide the data into 17,856 training examples and 1,985 validation examples. To test the generalizability of our approach across model scales, we fine-tune five open-weight models, namely Qwen2.5-0.5B, Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B, and Qwen2.5-14B on this corpus. The compute we train the models on are AMD Radeon Instinct MI250X GPUs and it took a total of 467 GPU hours.
Supervised Objective.
To fine-tune the models, formally, for each prompt and target sequence we minimize the token-level cross-entropy
@@ 4.1 Fine-Tuning Large Language Models
All models train for three epochs with a global batch size of 16 and context lengths of 131K for Qwen2.5 and 128K for the Llama variants. We use AdamW with , (Loshchilov and Hutter 2019), , weight decay , and gradient clipping at norm . The learning rate peaks at after a linear warm-up covering the first five percent of steps and then follows a linear decay.
@@
4.2 Multi-faceted Evaluation
Human Evaluation.
@@ 4.2 Multi-faceted Evaluation
After fine-tuning the language models with the assignment–feedback pairs, to test the performance of SEFL, we do human evaluation. We randomly sample 150 instances from the validation set. For each item, both the original instruction tuned model (A) and the model further fine-tuned with SEFL (B) produce feedback. Three human experts compared pairs of feedback responses produced for the same assignment and answer. For each item they read the original prompt, the student submission, and the two candidate feedback drafts from the non-tuned and SEFL-tuned model. Then, they select the feedback, from model A or B, that is better based on four base criteria:
@@
Accuracy. The generated feedback text refers to concrete strengths and weaknesses in the student work and avoid superficial remarks.
Actionability. Suggestions are clear, specific, and realistic for a student to apply.
Conciseness. Wording is brief and focused, with little repetition.
@@
Tone. Language stays constructive and professional while recognizing good elements.
@@
Raters were reminded to value efficiency over length, to prefer targeted advice over general principles, and to ignore formatting tricks. They recorded their choice as A or B and could leave an optional free-text comment. With this we calculate win rate (i.e., the percentage of choosing one feedback text over the other). This has become a de facto standard to evaluate long form text against each other (e.g., Rafailov et al. 2023).
@@
Each row took at most ten minutes, and the guidelines stressed taking regular breaks to sustain attention. We deliberately remove the tie option because a forced choice gives more informative labels and reduces hesitation, while a separate checkbox still lets raters mark assignmentanswerfeedback tuples as unrelated or nonsensical. This happened around 12% of the time, especially in the smaller, less capable models. The full annotation guidelines reported in the supplementary material in the code repository.
@@
Our human raters are in the aged 20–40 and from Europe. One identifies as female and the other two identifies as male. One female and male have a background in Computer Science and one male in Education Engineering. All have extensive experience in teaching and supervision or being taught and supervised, they all work in higher education (at different levels, i.e., research assistant and assistant professors) with near-native English proficiency.
LLM-as-a-Judge.

### v3

# SEFL: A Framework for Generating Synthetic Educational Assignment Feedback with LLM Agents

@@ Abstract
Providing high-quality feedback on student assignments is crucial for student success, but it is heavily limited by time and budgetary constraints.
In this work, we introduce Synthetic Educational Feedback Loops (SEFL), a synthetic data framework designed to generate data that resembles immediate, on-demand feedback at scale without relying on extensive, real-world student assignments and teacher feedback.
To obtain this type of data, two large language models (LLMs) operate in a teacher-student role to simulate assignment completion and formative feedback, generating 19.8K synthetic pairs of student work and corresponding critiques and actionable improvements from a teacher.
With this data, we fine-tune smaller, more computationally efficient LLMs on these synthetic pairs, enabling them to replicate key features of high-quality, goal-oriented feedback.
Through comprehensive evaluations with three LLM judges and three human experts, across a subset of 900 outputs, we demonstrate that SEFL-tuned models outperform both their untuned counterparts and an existing baseline in terms of feedback quality.
The potential for societal impact is reinforced by extensive qualitative comments and ratings from human stakeholders — both students and higher education instructors. SEFL has the potential to transform feedback processes for higher education and beyond.
Keywords: Synthetic Data, Education, Large Language Model, Agents
SEFL: A Framework for Generating Synthetic Educational Assignment Feedback with LLM Agents
| Mike Zhang, Amalie Pernille Dilling, Léon Gondelman |
| Niels Erik Ruan Lyngdorf, Euan D. Lindsay, Johannes Bjerva |
| University of Copenhagen, Denmark |
| Aalborg University, Denmark |
| Pioneer Centre for AI, Denmark |
| mike.zhang@di.ku.dk |
Abstract content
1. Introduction
@@
Constructive feedback is a cornerstone of higher education, promoting critical thinking and fostering deeper understanding Hattie (2008); Costello and Crane (2013). In higher education settings, however, providing consistent, high-quality feedback is complicated by privacy, consent, and transparency considerations in data collection Fischer et al. (2020); Suresh et al. (2022); Demszky and Hill (2023); Wang and Demszky (2024); Wang et al. (2024b); Lindsay et al. (2024), in addition to being a labor-intensive task. Advances in language technology offer opportunities to automate and augment higher education feedback processes, addressing these limitations.
@@
In particular, LLMs have shown progress in education Wang et al. (2024c), including automated grading Ke and Ng (2019); Ramesh and Sanampudi (2022); Stahl et al. (2024) and personalized tutoring Yun et al. (2024); Liu et al. (2024c); Rooein and Hovy (2024); Ross and Andreas (2024); Kwon et al. (2024); Zhang et al. (2024, 2025); Wang et al. (2024a). Yet, automating teacher-student assignment feedback with LLMs remains an open question. We seek to answer: RQ. How can synthetic teacher-student interactions generated by LLMs be leveraged to enable scalable and effective educational student assessment feedback?
@@
Here, we introduce Synthetic Educational Feedback Loops (SEFL), a framework that generates synthetic teacher-student interactions using LLM agents. In this framework, two LLMs, one acting as the teacher and the other as the student, simulate formative feedback workflows Conole and Oliver (2006); Nicol (2007). This synthetic data from the agents is then used to fine-tune smaller autoregressive language models, resulting in better feedback models that enable the development of scalable educational feedback systems. These systems can operate efficiently on the modest computational infrastructure available in higher education institutions, without requiring access to privacy-sensitive data.
@@
We show that: (i) Empirically, SEFL-tuned models outperform their non-tuned versions in win rate evaluations by three LLM judges and three human experts in giving assessment feedback. (ii) Through empirical and qualitative analysis of expert annotator comments, we find that larger models tend to provide more actionable, goal-oriented, and user-friendly feedback. They are also more consistent and better at supporting student autonomy. (iii) We observe strong agreement among human evaluators regarding feedback quality. (iv) We compare our approach to an existing baseline (Book2Dial; Wang et al., 2024a) and show that SEFL-tuned models provide better feedback according to all LLM judges.
@@
We contribute the following:
SEFL: An agentic framework that simulates teacher-student feedback loops with paired language model agents.
19,841 synthetic assignment–feedback pairs generated by SEFL to fine-tune smaller language models.
A comprehensive set of human and LLM evaluations that outline the strengths and limitations of SEFL with extensive qualitative analysis of the feedback provided by human experts.
Open-sourcing of all models, code, and data. 111Resources and Supplementary Material such as full prompts and example outputs can be found at https://github.com/jjzha/sefl and https://tinyurl.com/3zdu847k (HuggingFace).
2. Related Work
@@
Large language models are now supporting a broad spectrum of educational tasks. In automated grading, they score short answers, essays, and even programming assignments with accuracy that approaches expert instructors, which eases faculty workload and releases time for mentoring Ke and Ng (2019); Ramesh and Sanampudi (2022); Stahl et al. (2024). For personalized tutoring, conversational agents powered by these models adapt explanations, hints, and examples to each learner’s background knowledge and preferred style, producing measurable gains in engagement and achievement Yun et al. (2024); Liu et al. (2024c); Rooein and Hovy (2024); Ross and Andreas (2024); Kwon et al. (2024); Zhang et al. (2024). Research on peer learning shows that the same technology can mediate small-group discussions, suggest prompts, and highlight diverse viewpoints, leading to richer collaboration Bauer et al. (2023). In mathematics, aligning word problems and proofs to grade-level objectives has been automated, yielding encouraging results and reducing the manual effort required to curate question banks Botelho et al. (2023). Critical thinking curricula also benefit; LLMs can challenge students to justify claims, detect fallacies, and refine arguments in real-time Guerraoui et al. (2023). The models have also begun to assist scholars: studies report successful use for screening literature, summarizing drafts, and aligning reviewer comments with revision plans Liang et al. (2024); Sonkar et al. (2024). Complementing these functions are analytics tools that track learning trajectories and surface early warnings when a student slips behind Schwarz et al. (2018); Aslan et al. (2019); Alrajhi et al. (2021).
Despite this growing body of work, prior studies have rarely targeted the systematic generation of feedback on open-ended student submissions and assignments. In this study, we address this gap by using LLMs to generate extensive comment sets that teachers can accept as is or adapt. Decades of scholarship define effective feedback as goal-oriented, actionable, timely, user-friendly, and consistent while fostering self-evaluation Carless et al. (2011); Wiggins (2012). Feedback that is long and verbose can confuse learners. Therefore, concise wording is usually preferable, and comments that arrive soon after the original effort drive steady improvement Wiggins (2012). By producing immediate responses and tailoring suggestions to rubric criteria, LLM-based systems stand well-positioned to satisfy these guidelines while operating at the classroom and institution scale.
@@
Recent research shows how collaborative agentic LLMs can synthesize large-scale interactional datasets for educational tasks. For example, CAMEL Li et al. (2023) employs cooperative role-based dialogues to achieve shared objectives, whereas SimSeek Kim et al. (2022) utilizes agent-based conversations to construct comprehensive information-seeking datasets. In education, SocraticLM Liu et al. (2024b) simulates Socratic tutoring through multi-turn dialogue, and Book2Dial Wang et al. (2024a) generates teacher-student conversations from textbooks. In contrast, SEFL focuses on concise teacher-student feedback loops rather than extended instructional dialogues. While Nair et al. (2024) explore iterative revisions, SEFL generates diverse feedback pairs from assignment-answer-feedback tuples, enabling fine-tuning of smaller, cost-effective models for large-scale use.
3. SEFL: Synthetic Educational Feedback Loops
3.1. Synthetic Data Generation
@@
We employ a two-agent framework Wu et al. (2023) to simulate a student-teacher feedback loop, as seen in higher education. Both the teacher and student roles are simulated by two separate Llama-3.1-70B models for a two-turn conversation.222Note that if we mention a model, it is always the post-trained version (i.e., -Instruct). The models are tasked to generate assignmentanswerfeedback tuples. First, the student-agent requests an assignment using Fineweb-Edu Lozhkov et al. (2024) texts (Figure˜1), which is known for its educational content based on LLM judgments. Second, the teacher-agent creates an assignment that can be of any domain, e.g., STEM, social sciences, puzzles, and so forth. Then, the student-agent ( ) submits assignments containing several explicit errors, and the teacher-agent ( ) provides feedback. In and , we obtain an assignment-feedback pair. We deliberately avoided imposing strict rules on the types of errors that could be generated. We hypothesize that allowing the LLM to introduce errors naturally results in a more diverse and realistic set of mistakes than if we had a fixed checklist. We investigate both Qwen2.5-72B and Llama-3.1-70B for interactions. For a control check, we generate 5,000 interaction tuples with each model, and validate the output as an initial step to investigate the quality of initial feedback.
@@
We show in Table˜1 the results of this experiment. Out of 5,000 generated examples, Llama-3.1-70B generates 2,513 valid examples (i.e., valid JSON format and each feedback point refers to an error) compared to Qwen2.5-72B with 454 valid examples. For a further check, we use BERTScore Zhang et al. (2020) as a proxy to see whether each error-feedback pair of the valid generations relates to each other.333We only calculate it of the samples where both error and feedback have the same number of generations. We show that, regardless of Llama-3.1-70B generating more valid examples, the BERTScore (0.877) remains in a similar range to Qwen2.5-72B (0.919); in both cases, this indicates a high level of similarity. As a final qualitative check, we experimented with several prompts and consolidated them into a single, final prompt, which is included in the supplementary material. Finally, we use Llama-3.1-70B-generated data as the basis for all subsequent model fine-tuning, as it generated more valid examples. We spot-checked several prompts and consolidated the final full prompt in Figure˜6 (Appendix˜A).
@@
Data statistics.
After generating 5,000 examples, we continue to create example pairs, resulting in 19,841 teacher-student feedback pairs. In Table˜3, we present the final dataset statistics. We underline that the generation lengths for each agent are intentionally kept concise (170 subword tokens), based on the hypothesis that overly lengthy feedback may be counterproductive. This is in line with observations from Ferguson (2011), who finds that students tend to favor brief comments. We argue that balancing supportive and critical feedback is crucial, as LLMs often produce verbose responses by default, which can influence the preferences of both humans and LMs Saito et al. (2023).
Task Errors.
Once all data was generated, we took a random subsample of 226 instances with 428 explicit errors. We investigated what category of intentional errors the student agent generates. In total, we identified seven error types from the 428 errors and show the proportion in Table˜2. Most errors are of type “Omission and/or incompleteness” (surmounted for 50% of the errors):
1.
Omission and/or incompleteness: This includes missing details, lacking examples or evidence, not exploring a topic deeply enough, or failing to address all parts of the assignment.
2.
Stylistic and/or formatting issues: Problems with grammar, spelling, punctuation, tone (e.g., too informal), word count, and incorrect formatting (like improper citations).
3.
Factual inaccuracy: This includes incorrect dates, numbers, names, scientific facts, historical events, or misquoting a source text.
4.
Logical flaws and/or weak argumentation: These are errors in reasoning. They include making incorrect assumptions, oversimplifying complex topics, drawing faulty conclusions, or failing to construct a coherent and well-supported argument.
5.
Technical and/or procedural errors: This category is for mistakes in technical execution, such as incorrect mathematical calculations, flawed experimental design, or errors in code.
6.
Structural and organizational problems: These errors relate to the overall structure and flow of the response, such as a lack of clear transitions between paragraphs, poor organization, or a missing introduction/conclusion.
7.
Task misinterpretation: Student agent misunderstands the core requirement of the task, such as writing a story instead of an essay or failing to use a requested framework.
| Error Category | Count | % | Example Error Explanation from Student Agent |
| Omission and/or Incompleteness | 214 | 50.0% | “Failure to provide specific examples and evidence to support claims”. |
| Stylistic and/or Formatting Issues | 67 | 15.7% | “Incorrect spelling of the word ’protection’ in the last sentence of the answer”. |
| Factual Inaccuracy | 59 | 13.8% | “Incorrectly claimed Ibn al-Haytham was born in Egypt instead of Basra, Iraq”. |
| Logical Flaws and/or Weak Argumentation | 53 | 12.4% | “The statement about DSP chips having limitations is an oversimplification, as they are still widely used for various applications”. |
| Technical and/or Procedural Errors | 16 | 3.7% | “Incorrect calculation of the target waiting time, should be a 30% reduction from 10 minutes”. |
| Structural and/or Organizational Problems | 12 | 2.8% | “Lack of clear structure and organization in the comparative analysis”. |
| Task Misinterpretation | 7 | 1.6% | “The writing deploys facts that were not mentioned in the letter, like United States’ interest in containing the spread of communism”. |
| Total | 428 | 100.0% | — |
4. Methodology
4.1. Fine-Tuning Large Language Models
We divide the data into 17,856 training examples and 1,985 validation examples. To test the feedback quality of our approach across model scales, we fine-tune five open-weight models, namely Qwen2.5-0.5B, Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B, and Qwen2.5-14B on this synthetic dataset. The compute we train the models on is AMD Radeon Instinct MI250X GPUs, and it took a total of 467 GPU hours.
Training Objective.
To fine-tune the LLMs, for each prompt and target sequence we minimize the token-level cross-entropy
@@ 4.1. Fine-Tuning Large Language Models
All models train for three epochs with a global batch size of 16 and context lengths of 131K for Qwen2.5 and 128K for the Llama variants. We use AdamW with , Loshchilov and Hutter (2019), , weight decay , and gradient clipping at norm . The learning rate peaks at after a linear warm-up covering the first five percent of steps and then follows a linear decay.
@@
4.2. Multi-faceted Evaluation
4.2.1. Human Evaluation
@@
After fine-tuning the language models with the assignment-feedback pairs and testing the performance of SEFL, we conduct a human evaluation on a total of 450 instances with three annotators, following a similar approach to SoctraticLM Liu et al. (2024b), which annotated around 1,000 dialogues with 10 annotators. We randomly sample the subset of instances from the validation set. For each item, both the original instruction tuned model (A) and the model further fine-tuned with SEFL (B) produce feedback. Three human experts compared pairs of feedback responses produced for the same assignment and answer. For each item they read the original prompt, the student submission, and the two candidate feedback drafts from the non-tuned and SEFL-tuned model. Then, they select the feedback from model A or B that is better based on four base criteria:
@@
Accuracy: The generated feedback text focuses on concrete strengths and weaknesses in the student’s work, avoiding superficial remarks.
Actionability: Suggestions are clear, specific, and realistic for a student to apply.
Conciseness: Wording is brief and focused, with little repetition.
@@
Tone: Language stays constructive and professional while recognizing good elements.
@@
Raters were reminded to value efficiency over length, to prefer targeted advice over general principles, and to ignore formatting tricks. They recorded their choice as A or B and could leave an optional free-text comment. With this, we calculate the win rate (i.e., the percentage of choosing one feedback text over the other). This has become a de facto standard to evaluate long-form text against each other (e.g., Rafailov et al., 2023).
@@
Each row took at most ten minutes, and the guidelines stressed taking regular breaks to sustain attention. We deliberately remove the tie option because a forced choice provides more informative labels and reduces hesitation. At the same time, a separate checkbox still allows raters to mark assignment answer feedback tuples as unrelated or nonsensical. This happened around 12% of the time, especially in the smaller, less capable models. The full annotation guidelines are reported in the supplementary material (Table˜7; Appendix˜B).
Annotator Demographics.
@@
Our raters are aged 20-40 and from Europe. One identifies as female and the other two identify as male. One female and one male have a background in Computer Science, and one male has a background in Engineering Education. All have extensive experience in teaching and supervision, or being taught and supervised; they all work in higher education (at various levels, such as research assistant and assistant professors) with near-native English proficiency.
4.2.2. LLM-as-a-Judge